A FIELD SHAPE LIKE A TRIANGLE HAS A PRIMETER OF 224 METERS.THE LENGTH OF THE SECOND SIDE IS TWICE THE LENGTH OF THE FIRST SIDE,AND THE THIRD SIDE IS HALF THE LENGTH OF THE FIRST SIDE.FIND THE LENGTHS OF THE THREE SIDES.
Answer by SuperGirl98734(12)

if...
frist side = a
second side = 2a
third side = a/2
then...
a+2a+a/2=224
3a+a/2=224
(2)3a+a/2=224(2)
6a+a=448
7a=448
a=64
so...
first side = 64
second side = 128
third side = 32
i hope that helped and wasnt too hard to understand :)